# Change-Agent: Towards Interactive Comprehensive Change Interpretation and Analysis from Change Detection and Change Captioning

### v1

@@ Abstract
Monitoring changes in the Earth’s surface is crucial for understanding natural processes and human impacts, necessitating precise and comprehensive interpretation methodologies. Remote sensing satellite imagery offers a unique perspective for monitoring these changes, leading to the emergence of remote sensing image change interpretation (RSICI) as a significant research focus. Current RSICI technology encompasses change detection and change captioning, each with its limitations in providing comprehensive interpretation. To address this, we propose an interactive Change-Agent which integrates a multi-level change interpretation (MCI) model as eyes and a large language model (LLM) as the brain. Our Change-Agent can follow user instructions to achieve comprehensive change interpretation and insightful analysis according to user instructions, such as change detection and change captioning, change object counting, change cause analysis, etc. Our proposed MCI model contains two branches of pixel-level change detection and semantic-level change captioning, in which multiple BI-temporal Iterative Interaction (BI3) layers utilize Local Perception Enhancement (LPE) and the Global Difference Fusion Attention (GDFA) modules to enhance the model’s discriminative feature representation capabilities. To train the MCI model, we build the LEVIR-MCI dataset with change masks and captions of bi-temporal images. Extensive experiments demonstrate the effectiveness of the proposed change interpretation model and highlight the promising potential of our Change-Agent in facilitating comprehensive and intelligent interpretation of surface changes. We will make our dataset and codebase of the change interpretation model and Change-Agent publicly available to facilitate future research at https://github.com/Chen-Yang-Liu/Change-Agent
@@
Interactive Change-Agent, change captioning, change detection, multi-level change interpretation, large language model.
@@ I Introduction
Recent advancements in LLMs such as ChatGPT [16] and Llama2 [17] have been remarkable. Trained on vast corpora, LLMs have acquired extensive knowledge and demonstrated powerful capabilities in instruction comprehension, planning, reasoning, question answering, and text generation [18, 19]. These notable progress have provided robust support for our LLM-based Change-Agent. In this work, the LLM serves as the brain of the Change-Agent, assuming a pivotal role in internal scheduling, planing, and user-intent understanding. Integration with LLMs imbues our agent with enhanced flexibility and intelligence, enabling a nuanced understanding of user intentions and provision of customized change interpretation and intelligent analysis services such as changed object counting, estimation of change causes, prediction of future changes, etc.
@@ III-B Scale and Deformation of Changed Objects
In Fig. 4, we present an overview of the proposed Change-Agent. Comprising a MCI model as eyes and an LLM as the brain, the Change-Agent is endowed with visual perception and cognitive capabilities. This integration enables the agent to achieve an interactive and comprehensive interpretation and analysis of surface changes.
a comprehensive understanding and interpretation of surface changes, empowering the agent to provide insightful analyses and decision-making support. The MCI model facilitates robust and comprehensive change interpretation capabilities, while the LLM orchestrates the planning and information processing aspects of the agent. In the subsequent sections, we delve into the construction and training of the multi-level change interpretation model and elucidate the LLM-based scheduling of the Change-Agent.
@@ V-C Experimental Details
We implemented all models using the PyTorch deep learning framework and trained them on the NVIDIA GTX 4090 GPU. During training, we minimized the loss function defined by Equation 20 and employed the Adam optimizer [71] with an initial learning rate of 0.0001 to optimize the model parameters. We set the maximum number of epochs to 200 and utilized word embeddings with a dimensionality of 512. The backbone for feature extraction is the Siamese weight-shared Segformer-B1 [52]. Initially, we trained the backbone until the sum of BLEU and MIoU indicators did not increase for 50 consecutive epochs. Subsequently, we froze the backbone network and continued training the two branches separately.
@@ V-D Multi-level Change Interpretation Performance
Method MIoU BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGEL CIDEr-D
@@
  +GDFA 86.41 85.32 77.45 70.63 65.10 39.59 74.45 135.94
    +GDFA+LPE 86.43 86.68 78.75 71.74 65.95 40.80 75.96 140.29
@@
V-E Ablation Studies
V-E1 BI3 Layer
As a critical component of our change interpretation model, the BI3 layer incorporates LPE and GDFA modules. The LPE module enhances feature representation through local perception enhancement,improving the model’s sensitivity to multi-scale changes. The GDFA module effectively integrates the differential information, allowing the model to capture the change area more accurately. In Table III, we conducted ablation experiments on the effectiveness of the LPE and GDFA modules. The baseline utilizes Transformer encoding layers to process bi-temporal features. The experimental results show that integrating the LPE and GDFA modules leads to improved performance in both change detection and change captioning tasks.
@@ V-E2 Balancing Change Detection and Captioning
In our exploration of multi-task learning for change detection and change captioning, achieving an effective balance between these tasks posed a significant challenge. Initially, when we simply combined the individual losses $\mathcal{L}{det}$ and $\mathcal{L}{cap}$ to form the total loss and minimized it to optimize the multi-task model parameters, we observed large discrepancies in the magnitudes of these two losses, as illustrated in Fig. 10. This imbalance during training could potentially impact overall change interpretation performance. To address this issue, we adopt a balancing strategy detailed in Equation 20 to normalize the two losses and achieve the training balance between the two tasks. Table IV shows the impact of these two strategies on model performance. It’s evident that our balancing strategy plays a crucial role in harmonizing the training process, leading to more balanced performance across both change detection and captioning tasks.
@@ V-F Qualitative Change Interpretation Results
To illustrate the effectiveness of our proposed model in simultaneously generating change masks and change descriptions, Fig. 11 presents comparative results between our proposed model and the baseline model. The experimental results demonstrate the superior capabilities of our model in both change detection and captioning.
In terms of change detection, our model exhibits remarkable proficiency in identifying subtle changes in small buildings, as depicted in the first row of Fig. 11. Furthermore, our model can produce more refined road masks in rows 5 and 7 compared to the baseline. In contrast to the baseline, which erroneously detects some buildings in the sixth and eighth rows, our model performs admirably.
Regarding change captioning, in the fourth and eighth rows of Fig. 11, our model can accurately recognize and describe the changes in the bi-temporal images. Conversely, the baseline erroneously concludes that no changes have occurred, resulting in a critical error in the description. Furthermore, it is evident from the first four rows of Fig. 11 that our model can not only accurately identify the types of changing objects but also describe their spatial positions and the relationships between the changing objects and surrounding features.
Figure 12: Examples of conversations between a user and our Change-Agent. Our Change-Agent can understand users and carry out tasks such as change mask prediction, change captioning, change object counting, mask post-processing, change causes estimation, and future change prediction.
Figure 13: The executable Python function code generated by LLM when a user requests to only detect changed buildings and display it in red.
@@ V-G Interaction Between User and Change-Agent
Leveraging the change interpretation model trained on the LEVIR-MIC dataset, we have constructed an LLM-based Change-Agent. The change interpretation model provides our agent with powerful change interpretation capability. The LLM enables our agents to understand user instructions and call tools to achieve interactive and comprehensive change interpretation and processing.
In Fig. 12, we illustrate two examples of conversations between a user and our Change-Agent. In the first conversation, our agent showcases its adeptness in understanding user requirements and subsequently carries out tasks such as change mask prediction, change captioning, change object counting, and mask post-processing. This highly automated processing greatly enhances user efficiency and provides users with a more convenient and intuitive experience in change interpretation. In the second conversation, our agent can describe the changes in the images with the help of change interpretation model, and use the rich knowledge inherent in the LLM to answer the user’s questions about the cause of the changes and the prediction of future changes. This will assist user decision-making.
Fig. 13 presents a specific example to illustrate the practical application value of our Change-Agent. When a user requests our agent to only detect changed building masks and display building areas in red, the LLM within our agent can accurately understand this instruction and generate executable Python function code.
Through these interactive capabilities, our Change-Agent serves as a versatile tool for facilitating nuanced and insightful analysis of changes within remote sensing images, fostering seamless communication and collaboration between users and remote sensing expertise.
@@ VI Discussion and Future
Our research introduces the LEVIR-MIC dataset that lays a valuable data foundation for facilitating multi-task learning of change detection and change captioning and exploring the interconnection between two tasks. Besides, our Change-Agent opens up new avenues for change interpretation and brought a new thinking direction to the remote sensing field. However, there are still some issues worthy of in-depth exploration. Future work can be carried out from the following aspects:
  • Multi-Task Learning. Based on the proposed dataset, we initially explored multi-task learning of change detection and change captioning and found some challenges, such as the balance of two tasks. The intrinsic relationship between the two tasks still needs to be deeply explored. By designing better model structures and training strategies, future work can focus on further promoting the mutual promotion between the two tasks, thereby improving the overall interpretation performance.
  • Scheduling of Change-Agent. The system prompts designed by humans and inputted into the LLM play a crucial role in facilitating accurate agent scheduling. By optimizing the prompt design, we can enable the agent to more accurately capture the user’s intentions and plan a more reasonable and efficient task execution pathway.
  • Tool Expansion. Beyond change interpretation, we can provide the agent with additional models to enhance its capabilities to handle more complex and comprehensive remote sensing image processing tasks, thereby providing users with more comprehensive and professional interpretation services. Additionally, the updateability of tools in the agent allows us to introduce new tools and models, ensuring that our agent remain up-to-date.
  • Multi-Agent Systems. By introducing multiple agents with diverse capabilities and fostering collaboration among them, researchers can achieve more flexible and collaborative image interpretation work. It is helpful for building a more efficient and intelligent remote sensing system.
@@ VII Conclusion
In this paper, we have addressed the critical need for comprehensive and intelligent interpretation of surface changes on Earth through the development of a novel Change-Agent. Our agent is propelled by a multi-level change interpretation model and an LLM. The change interpretation model endows the agent with robust visual change perception capabilities, while the LLM orchestrates the agent’s scheduling. The proposed change interpretation model encompasses dual branches of change detection and captioning, capable of providing both pixel-level change masks and semantic-level understanding in textual form. Within the dual-branch framework, we introduce a BI3 layer with LPE and GDFA to enhance the model’s discriminative feature representation capabilities. Additionally, we build a dataset comprising diverse change detection masks and descriptions to support model training. Experiments validate the efficacy of the proposed change interpretation model and highlight the promising potential of our Change-Agent in facilitating comprehensive and intelligent interpretation of surface changes.

### v2

# Change-Agent: Towards Interactive Comprehensive Remote Sensing Change Interpretation and Analysis

@@ Abstract
Monitoring changes in the Earth’s surface is crucial for understanding natural processes and human impacts, necessitating precise and comprehensive interpretation methodologies. Remote sensing satellite imagery offers a unique perspective for monitoring these changes, leading to the emergence of remote sensing image change interpretation (RSICI) as a significant research focus. Current RSICI technology encompasses change detection and change captioning, each with its limitations in providing comprehensive interpretation. To address this, we propose an interactive Change-Agent, which can follow user instructions to achieve comprehensive change interpretation and insightful analysis according to user instructions, such as change detection and change captioning, change object counting, change cause analysis, etc. The Change-Agent integrates a multi-level change interpretation (MCI) model as the eyes and a large language model (LLM) as the brain. The MCI model contains two branches of pixel-level change detection and semantic-level change captioning, in which multiple BI-temporal Iterative Interaction (BI3) layers utilize Local Perception Enhancement (LPE) and the Global Difference Fusion Attention (GDFA) modules to enhance the model’s discriminative feature representation capabilities. To support the training of the MCI model, we build the LEVIR-MCI dataset with a large number of change masks and captions of changes. Extensive experiments demonstrate the effectiveness of the proposed MCI model and highlight the promising potential of our Change-Agent in facilitating comprehensive and intelligent interpretation of surface changes. To facilitate future research, we will make our dataset and codebase of the MCI model and Change-Agent publicly available at https://github.com/Chen-Yang-Liu/Change-Agent
@@
Interactive Change-Agent, change captioning, change detection, multi-task learning, large language model.
@@ I Introduction
Recent advancements in LLMs such as ChatGPT [16] and Llama2 [17] have been remarkable. Trained on vast corpora, LLMs have acquired extensive knowledge and demonstrated powerful capabilities in instruction comprehension, planning, reasoning, question answering, and text generation [18, 19]. These notable progress have provided robust support for our LLM-based Change-Agent. In this work, the LLM serves as the brain of the Change-Agent, assuming a pivotal role in internal scheduling, planning, and user-intent understanding. Integration with LLMs imbues our agent with enhanced flexibility and intelligence, enabling a nuanced understanding of user intentions and provision of customized change interpretation and intelligent analysis services such as changed object counting, estimation of change causes, prediction of future changes, etc.
@@ V-C Experimental Details
We implemented all models using the PyTorch deep learning framework and trained them on the NVIDIA GTX 4090 GPU. During training, we minimized the loss function defined by Equation 20 and employed the Adam optimizer [71] with an initial learning rate of 0.0001 to optimize the model parameters. We set the maximum number of epochs to 200 and utilized word embeddings with a dimensionality of 512. The backbone for feature extraction is the Siamese weight-shared Segformer-B1 [52]. Initially, we trained the backbone until the sum of BLEU-4 and MIoU indicators did not increase for 50 consecutive epochs. Subsequently, we froze the backbone network and continued training the two branches separately.
@@ V-D Multi-level Change Interpretation Performance
Method GDFA LPE MIoU BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGEL CIDEr-D
@@
MCI-G 86.41 85.32 77.45 70.63 65.10 39.59 74.45 135.94
MCI-L 86.42 86.43 78.53 71.54 65.75 40.45 75.72 138.99
Our MCI 86.43 86.68 78.75 71.74 65.95 40.80 75.96 140.29
V-E Ablation Studies
V-E1 BI3 Layer
As a critical component of our MCI model, the BI3 layer incorporates LPE and GDFA modules. The LPE module enhances feature representation through local perception enhancement,improving the model’s sensitivity to multi-scale changes. The GDFA module effectively integrates the differential information, allowing the model to capture the change area more accurately. In Table III, we conducted ablation experiments on the effectiveness of the LPE and GDFA modules. The baseline utilizes Transformer encoding layers to process bi-temporal features. The experimental results show that integrating the LPE and GDFA modules leads to improved performance in both change detection and change captioning tasks.
@@ V-E2 Balancing Change Detection and Captioning
In our exploration of multi-task learning for change detection and change captioning, achieving an effective balance between these tasks posed a significant challenge. Initially, when we simply combined the individual losses $\mathcal{L}_{det}$ and $\mathcal{L}_{cap}$ to form the total loss and minimized it to optimize the multi-task model parameters, we observed large discrepancies in the magnitudes of these two losses, as illustrated in Fig. 10. This imbalance during training could potentially impact overall change interpretation performance. To address this issue, we adopt a balancing strategy detailed in Equation 20 to normalize the two losses and achieve the training balance between the two tasks. Table IV shows the impact of these two strategies on model performance. It’s evident that our balancing strategy plays a crucial role in harmonizing the training process, leading to more balanced performance across both change detection and captioning tasks.
@@ V-F Qualitative Change Interpretation Results
To illustrate the effectiveness of our proposed model in simultaneously generating change masks and change descriptions, Fig. 11 presents comparative results between our proposed model and the baseline model. The experimental results demonstrate the superior capabilities of our model in joint interpretation of change detection and change captioning.
In terms of change detection, our model exhibits remarkable proficiency in identifying subtle changes in small buildings, as depicted in the first row of Fig. 11. Moreover, our model produces more refined road masks in rows 5 and 7 compared to the baseline. Unlike the baseline, which erroneously detects some buildings in the sixth and eighth rows, our model performs admirably.
Regarding change captioning, our model accurately recognizes and describes the changes in the bi-temporal images in the fourth and eighth rows of Fig. 11. Conversely, the baseline erroneously concludes that no changes have occurred, resulting in a critical error in the description. Furthermore, it is evident from the first four rows of Fig. 11 that our model not only accurately identifies the types of changing objects but also describes their spatial positions and the relationships between the changing objects and surrounding features.
Figure 12: Conversations between users and our Change-Agent. Our Change-Agent can understand users and carry out tasks such as change mask prediction, change captioning, change object counting, mask post-processing, change causes estimation, and future change prediction.
Figure 13: The executable Python function code generated by LLM within our Change-Agent when a user requests to perform change detection, display building areas in green, display road areas in blue, and count changed buildings.
@@ V-G Interaction Between User and Change-Agent
Utilizing the MCI model trained on the LEVIR-MIC dataset, we have constructed an LLM-based Change-Agent. The MCI model provides our agent with powerful change interpretation capability. The LLM serves as the brain behind our agent, facilitating seamless interaction and comprehensive interpretation of changes.
In Fig. 12, we provide two examples of conversations between users and our Change-Agent, illustrating its adaptability and versatility. In the first conversation, our agent showcases its adeptness in understanding user instructions and subsequently carries out tasks such as change mask prediction, change captioning, changed object counting, and mask post-processing. This automated approach significantly enhances user efficiency and offers a more intuitive experience in change interpretation. In the second conversation, our agent leverages the MCI model to describe the changes in the images, and leverages the rich knowledge inherent in the LLM to provide insightful answers to user queries regarding change causality and future change predictions, aiding in user’s decision-making processes.
Fig. 13 presents a specific example to illustrate the executable code generated by the LLM within our Change-Agent. When the user requests to perform change detection, display building areas in green, display road areas in blue, and count changed buildings, the LLM efficiently generates executable Python code, demonstrating the agent’s versatility and responsiveness to user needs.
Through these interactive capabilities, our Change-Agent serves as a versatile tool for facilitating nuanced and insightful analysis of changes in remote sensing images, fostering seamless communication and collaboration between users and remote sensing expertise.
@@ VI Discussion and Future
Our research introduces the LEVIR-MIC dataset, which lays a solid data foundation for facilitating multi-task learning and exploring the interconnection between change detection and change captioning. Besides, our Change-Agent opens up new avenues for change interpretation and brought a new thinking to the remote sensing field. However, there are still some issues worthy of further exploration:
  • Multi-Task Learning. Challenges such as task balance and inter-task relationships remain in multi-task learning. Future work can focus on designing better model structures and training strategies to improve overall interpretation performance on change detection and change captioning.
  • Scheduling of Change-Agent. The hand-craft system prompts inputted into the LLM play a crucial role in facilitating accurate agent scheduling. Optimizing system prompts could enable to better grasp user intentions and plan more accurate and rational task execution pathways.
  • Tool Expansion. Beyond change interpretation, providing the agent with additional models can enhance its capabilities for handling complex remote sensing image processing tasks, offering users more comprehensive interpretation services. Additionally, the agent’s updateability ensures it remains up-to-date with new tools and models.
  • Multi-Agent Systems. Introducing multiple agents with diverse capabilities and fostering collaboration among them could lead to more flexible and collaborative image processing and analysis, contributing to the development of more efficient and intelligent remote sensing systems.
@@ VII Conclusion
In this paper, we have addressed the critical need for a comprehensive and intelligent interpretation of Earth’s surface changes through developing a novel Change-Agent. Our Change-Agent can follow user instructions to achieve comprehensive change interpretation and insightful analysis according to user instructions. The Change-Agent, fueled by an MCI model and an LLM, exhibits robust visual change perception capabilities and adept scheduling. The MCI model, comprising two branches of change detection and captioning, offers pixel-level change masks and semantic-level textual change descriptions. In these two branches, we propose BI3 layers with LPE and GDFA to enhance the model’s discriminative feature representation capabilities. Additionally, we build a dataset comprising diverse change detection masks and descriptions to support model training. Experiments validate the effectiveness of the proposed MCI model and underscore the promising potential of our Change-Agent in facilitating comprehensive and intelligent interpretation of surface changes.